If $7,000.00 is invested in a bank that pays 5% interest, how much more must be invested in bonds at 9% to earn 7% on the total investment?
To earn 7% on the total investment, $______ more must be invested in bonds at 9%.

Answer by richwmiller(17219)

one investment all 3 rates
0.05*7000+0.09(x)=0.07(7000+x)
Multiply out.
350+0.09(x)=490+0.07x
combine like terms
0.02x=140
isolate x
x=140/0.02
Divide
x=7000
check
0.05*7000+0.09(x)=0.07(7000+x)
350+0.09*7000=0.07(7000+7000)
350+630=0.07(14000)
980=980
ok